Grade 12Physical Chemistry

A golf ball has a mass of 40g, and a speed of 45 m/s. If the speed can be measured within accuracy of 2%, calculate the uncertainty in the position.

sudhanshu
12 Years agoGrade 12

3 Answers

Pankaj
11 Years ago
Given, The uncertainty in the speed is 2%
Therefore,
45 x 2/100 = 0.9ms-1 .
Now, dxdp = h/4\pi = (6.626 x 10-34 Js) / (4x3.14 x 40gx10-3 kg g-1 x 0.9m s-1)
=1.46 x10-33 m
Hence, the uncertainty in position is 1.46 x 10-33 m
irfan noushad
7 Years ago
Use Heisenberg's uncertainty principle, 
e.g., ∆x = h/4πm∆v
Here, ∆x is the uncertainty in position
∆v is the uncertainty in velocity
m is the mass of Particle .
Given, m = 40g = 0.04kg 
∆v = 2% of v = 2 × 45/100 = 0.9 m/s 
h = 6.626 × 10⁻³⁴ J.s
Now, ∆x = 6.626 × 10⁻³⁴/(4 × 3.14 × 0.04 × 0.9)
= 14.654 × 10⁻³⁴ m 
Hence, uncertainty in position = 1.4654 × 10⁻³³ m
Kushagra Madhukar
6 Years ago
Dear student,
 
As given in the question,
Δv = 2% of v = 2/100 x 45 = 0.9ms-1 .
Now, Δxp = h/4\pi = (6.626 x 10-34 Js) / (4x3.14 x 40gx10-3 kg g-1 x 0.9m s-1) = 1.46 x10-33 m
Hence, the uncertainty in position is 1.46 x 10-33 m
Hope it helps.
 
Thanks and regards,
Kushagra